# Aligning Multiple Knowledge Graphs in a Single Pass

### v1

@@ Abstract
Entity alignment (EA) is to identify equivalent entities across different knowledge graphs (KGs), which can help fuse these KGs into a more comprehensive one. Previous EA methods mainly focus on aligning a pair of KGs, and to the best of our knowledge, no existing EA method considers aligning multiple (more than two) KGs. To fill this research gap, in this work, we study a novel problem of aligning multiple KGs and propose an effective framework named MultiEA to solve the problem. First, we embed the entities of all the candidate KGs into a common feature space by a shared KG encoder. Then, we explore three alignment strategies to minimize the distances among pre-aligned entities. In particular, we propose an innovative inference enhancement technique to improve the alignment performance by incorporating high-order similarities. Finally, to verify the effectiveness of MultiEA, we construct two new real-world benchmark datasets and conduct extensive experiments on them. The results show that our MultiEA can effectively and efficiently align multiple KGs in a single pass.
@@
1 Introduction
knowledge graphs (KGs) are a special kind of graph that can store a wealth of structural facts (i.e. knowledge) about the real world. Each fact is usually structured as a triple $(h,r,t)$, representing that head (subject) entity $h$ and tail (object) entity $t$ hold relation $r$ between them. In recent years, KGs have successfully supported many web applications such as search engines [1], question-answer systems [2, 3, 4], recommender systems [5, 6, 4], etc.
In practice, different KGs are constructed based on diverse data sources and different extraction approaches, and a single KG can usually cover only a specific aspect of structural facts. For example, an English KG usually contains more facts about the English-speaking society, while a Chinese KG contains more facts about the Chinese-speaking society. Considering that more KGs together can provide more comprehensive structural facts from various aspects, researchers have proposed many methods (Cf. surveys [7, 8]) to fuse a pair of KGs into a unified one. The general process is to first identify equivalent entities between the candidate KGs, and then let them serve as the “bridge entities” to link the candidate KGs into a unified one.
@@ 1 Introduction
However, all the existing EA methods are specifically designed for aligning only a pair of candidate KGs, which can cover limited knowledge for downstream applications. To the best of our knowledge, none of the existing EA methods consider aligning multiple (more than two) KGs. For example, Figure 1 shows a case where four candidate KGs need to be aligned, which triggers a more challenging research problem.
There are three obstacles that prevent traditional pair-wise KG alignment methods from being applied to the multiple KG alignment problem studied in this work. First, existing EA datasets are particularly constructed for pair-wise KG alignment. At present, there is no benchmark dataset that can support the study of aligning multiple KGs. Second, if we trivially adapt existing methods to this task, they need to be separately executed multiple times since they can only utilize the local (pair-wise) alignment information in a single pass, which is inefficient. Last and most importantly, they cannot capture some useful global (beyond pair-wise) alignment information, affecting the final alignment performance. For the example in Figure 1, when they align KG-en with KG-fr, and align KG-en with KG-zh independently, the entities of the three KGs will be projected into four separate embedding spaces. Consequently, it is easy to yield inconsistent results like $e_{1}^{(en)}\equiv e_{1}^{(fr)}$, $e_{1}^{(en)}\equiv e_{1}^{(zh)}$, and $e_{1}^{(fr)}\equiv e_{2}^{(zh)}$ (wrong), violating the transitivity of the alignment relationship.
To fill the research gap in this field, in this work, we innovatively study the problem of aligning multiple KGs in a single pass. First of all, to facilitate the study of this problem, we construct two new benchmark datasets that contain multiple KGs to be aligned. Then, we formulate a novel research problem of aligning multiple KGs concurrently. Finally, we propose an effective framework named MultiEA to solve the problem. As illustrated in Figure 1, our MultiEA can concurrently align the four candidate KGs in a single pass, which is more practical. In addition, it embeds the entities of all the candidate KGs into a unified space and thus can capture useful global alignment information, which helps achieve better alignment performance. Recall the example above, if MultiEA pulls $e_{1}^{(en)}$ and $e_{1}^{(fr)}$ closer and pulls $e_{1}^{(en)}$ and $e_{1}^{(zh)}$ closer. Then, $e_{1}^{(fr)}$ and $e_{1}^{(zh)}$ will definitely be pulled closer as well, leading to the consistent result: $e_{1}^{(en)}\equiv e_{1}^{(fr)}$, $e_{1}^{(en)}\equiv e_{1}^{(zh)}$, and $e_{1}^{(fr)}\equiv e_{1}^{(zh)}$ (right).
@@
  • We conduct extensive experiments on the two constructed benchmark datasets. The results demonstrate that MultiEA can effectively and efficiently align multiple KGs in a single pass. We will release our source codes and datasets to facilitate further research on this problem.
2 Related Work
@@
Trans-based EA Methods. Early EA methods exploit KG embedding methods, e.g., TransE [15], to project a pair of candidate KGs into a common low-dimensional Euclidean space by letting the embeddings of head entity $h$, relation $r$, and tail entity $h$ satisfy the triangular relationship: $h+r\approx t$. Given a set of pre-aligned entity pairs (i.e., seed alignment labels), they minimize the distance between these equivalent entity pairs in the embedding space. Thus, more potential equivalent entity pairs can be discovered after model optimization. Representative Trans-based EA methods include JEwP [9], JAPE [12], MTransE [11], IPTransE [10], BootEA [13], TransEdge [31], MultiKE [14], etc.
GNN-based EA Methods. Motivated by the extraordinary success of graph neural networks (GNNs) in extracting the structural features of graphs, GCN-Align [16] first introduces GCN [30] into EA, which has shown superior performance than Trans-based EA Methods, e.g., JEwP [9], MTransE [11] and JAPE [12]. Further, AVR-GCN [17] deftly combines the graph convolutional operator in GNNs and the translation operator in Trans-based methods to extract better entity embeddings. Afterward, a series of following GNN-based EA methods, e.g.,  [18, 19, 20, 21, 22, 23, 26, 27, 29, 28, 24, 25, 32, 33, 34] are proposed. Recently, RREA [24] insightfully summarizes KG embedding-based EA methods and GNN-based EA methods into a unified framework consisting of two modules: shape-builder and alignment. The former constraints KGs into a specific distribution in the embedding space, and the latter minimizes the distance between the embeddings of pre-aligned entity pairs.
@@ 3 Preliminaries
As shown in Figure 1, KG-en, KG-zh, KG-fr, and KG-ja are four candidate KGs that need to be aligned, where the four entities marked with red dashed lines refer to the same object, i.e., “Sweden”. We treat the four equivalent entities as one of the seed alignment labels, to help our EA algorithm find more potential equivalent entities. Finally, we can link (merge) the four KGs based on these detected equivalent entities across the four KGs.
4 Framework
@@ 4.2 Training for Alignment
In this work, we propose three strategies for minimizing the distance among equivalent entities. The ideas are intuitively shown in Figure 2. For a positive example $l=(e^{(1)},e^{(2)},...,e^{(m)},...,e^{(M)})\in\mathcal{S}$, we firstly obtain its associated entity embeddings from the KG encoder and denote them as: $(\mathbf{h}^{(1)},\mathbf{h}^{(2)},...,\mathbf{h}^{(m)},...,\mathbf{h}^{(M)})$. Then, the distance $d(l)$ is minimized by three strategies as follows.
(1) Moving toward Mean. This strategy lets all the equivalent entities approach their mean, as illustrated in Figure 2(a). We first compute the mean as follows:
@@
(2) Moving toward Anchor. In practice, the candidate KGs are usually unbalanced due to various realistic factors. For instance, on our constructed multi-lingual KG dataset DBP-4, the English KG contains many more triples than the other three KGs (Cf. Table I). Therefore, as shown in Figure 2(b), we designate entities from KG DBP-4: en as anchor entities and let the equivalent entities from the other KGs approach these anchor entities. This idea is formally described as follows:
@@
(3) Moving toward Each Other. This strategy lets equivalent entities approach each other. As illustrated in Figure 2(c), we minimize the distances between all possible pairs of equivalent entities, formally described as follows:
@@
Input : $M$ KGs $\{\mathcal{G}^{(1)},\mathcal{G}^{(2)},...,\mathcal{G}^{(m)},...,\mathcal{G}^{(M)}\}$,
seed alignment labels $\mathcal{S}$.
1 Randomly initialize model parameters;
2 Select one distance metric function from Eqs. (11, 12), Eq. (13), or Eq. (14);
3 Augment KG data by Eqs. (1, 2);
4 Get neighbor relation-entity tuples by Eq. (3);
5 while not converge do
6      Compute relation-specific projection matrices by Eq. (5);
7      Compute attention coefficients by Eqs. (7, 8);
8      Perform multiple layers of aggregation by Eq. (4);
9      Randomly sample negative examples by Eq. (10);
10      Compute loss by Eq. (9);
11      Update model parameters by gradient descent;
12
13      end while
@@ 11      Update model parameters by gradient descent;
4.3 Training Time Complexity
@@
In the KG encoding phase (Cf. Section 4.1), the time complexity involves three main computational operations. First, the aggregation operation described by Eq. (4) has time complexity ${\mathcal{O}}(|\mathcal{E}|\cdot d^{2})$. Second, the computation of relation-specific projection matrices described by Eq. (5) has time complexity ${\mathcal{O}}(|\mathcal{R}|\cdot d^{2})$. Finally, as described by Eqs. (7, 8), the computation of attention coefficients has time complexity ${\mathcal{O}}(|\mathcal{T}|\cdot(d+d+d^{2}))$. Considering that in practice, the embedding dimensionality $d$ is a relatively small value, the time complexity of KG encoding is ${\mathcal{O}}(|\mathcal{E}|+|\mathcal{R}|+|\mathcal{T}|)$.
In the alignment phase (Cf. Section 4.2), as described by Eq. (12) and Eq. (13), both the mean strategy and the anchor strategy require time complexity ${\mathcal{O}}(M)$. As described by Eq. (14), the each other strategy requires time complexity ${\mathcal{O}}(M^{2})$. As described by Eq. (9), taking the number of labels into consideration. The mean strategy and the anchor strategy have time complexity of ${\mathcal{O}}(|\mathcal{S}|\cdot\eta\cdot M^{2})$, and each other strategy has time complexity ${\mathcal{O}}(|\mathcal{S}|\cdot\eta\cdot M^{3})$.
@@ 5.1 Benchmark Datasets
TABLE II: Comparison of multiple KG alignment accuracy (%). The best results are bolded, and the second best results are underlined. The notation denotes that we re-implement the baseline methods to make them compatible with the task.
Methods DBP-4 DWY-3
MTransE 3.40 19.72 32.64 25.83 55.94 65.31
IPTransE 4.11 23.36 39.55 28.74 60.02 68.84
GCN-Align 4.48 27.62 44.94 32.46 63.18 73.91
KECG 5.62 33.35 49.61 34.08 67.51 77.52
NAEA 6.14 34.13 48.07 35.17 69.88 79.43
PSR 7.84 38.17 50.97 37.55 71.85 80.13
MultiEA+mean-infer 5.95 40.48 51.02 36.68 69.33 79.65
MultiEA+anchor-infer 6.32 40.33 52.08 37.54 70.21 79.19
MultiEA+each-infer 6.17 41.70 52.0 36.76 71.52 79.18
MultiEA+mean+infer 10.03 48.18 59.89 40.54 73.90 81.24
MultiEA+anchor+infer 10.58 48.80 59.95 42.33 74.13 83.54
MultiEA+each+infer 10.25 50.38 61.75 41.19 75.44 85.42
5.2 Evaluation Metric
@@
5.3 Baselines and Variants
@@
Recall that in Section 4.2, we introduce three different alignment strategies. Here, we use +mean, +anchor, and +each to denote the three strategies, respectively. In Section 4.4, we propose the inference enhancement module. Here, we use +infer or -infer to indicate whether the module is equipped. By considering all the permutations, we set the following six variants for our MultiEA:
@@
Figure 5: Visualization of the embeddings of the ten groups of equivalent entities listed in Figure 3 and Figure 4.
5.4 Implementation Details
@@ 5.5 Multiple KG Alignment
5.6 Case Study
To intuitively show the effectiveness of our MultiEA, in Figure 3 and Figure 4, we list ten groups of equivalent entities discovered by MultiEA (the best-performing variant MultiEA+each+infer) on the two datasets. As we can see, MultiEA can effectively discover equivalent entities across all the candidate KGs.
Further, we visualize the embeddings of these entities. Specifically, we utilize the t-SNE algorithm [52] to project them into the 2-dimensional Euclidean space. Figure 5(a) and Figure 5(b) show the results on DBP-4 and DWY-3, respectively, where different colors and different shapes mark the entities from different KGs, and the equivalent entities are connected by black lines. As we can see, on both datasets, the ten groups of equivalent entities are obviously gathered together, indicating the effectiveness of the alignment mechanism of MultiEA. On the other hand, a proper margin is maintained between different groups, avoiding the risk of model overfitting. The result on DWY-3 shows a relatively better distribution. This is consistent with the higher accuracy on DWY-3 as reported in Table II.
5.7 Efficiency Study
In this experiment, we demonstrate the efficiency advantage of our MultiEA in aligning multiple KGs. As shown in Table III and Table IV, we first report the experimental results of our re-implemented GNN-based EA baselines as described in Section 5.5. That is, they align multiple candidate KGs by splitting this task into multiple separate pair-wise sub-tasks, as marked by in the tables. Then, we apply the “each other” alignment strategy and the inference enhancement module of our MultiEA+each+infer to each baseline X, as denoted as MultiEA (X) in the tables.

### v2

@@ Abstract
Entity alignment (EA) is to identify equivalent entities across different knowledge graphs (KGs), which can help fuse these KGs into a more comprehensive one. Previous EA methods mainly focus on aligning a pair of KGs, and to the best of our knowledge, no existing EA method considers aligning multiple (more than two) KGs. To fill this research gap, in this work, we study a novel problem of aligning multiple KGs and propose an effective framework named MultiEA to solve the problem. First, we embed the entities of all the candidate KGs into a common feature space by a shared KG encoder. Then, we explore three alignment strategies to minimize the distances among pre-aligned entities. In particular, we propose an innovative inference enhancement technique to improve the alignment performance by incorporating high-order similarities. Finally, to verify the effectiveness of MultiEA, we construct two new real-world benchmark datasets and conduct extensive experiments on them. The results show that our MultiEA can effectively and efficiently align multiple KGs in a single pass. We release the source codes of MultiEA at: https://github.com/kepsail/MultiEA.
@@
I Introduction
Knowledge graphs (KGs) are a special kind of graph that can store a wealth of structural facts (i.e. knowledge) about the real world. Each fact is usually structured as a triple $(h,r,t)$, representing that head (subject) entity $h$ and tail (object) entity $t$ hold relation $r$ between them. In recent years, KGs have successfully supported many web applications such as search engines [1], question-answer systems [2, 3, 4, 5], recommender systems [6, 7, 4], knowledge graph reasoning [8, 9], etc.
In practice, different KGs are constructed based on diverse data sources and different extraction approaches, and a single KG can usually cover only a specific aspect of structural facts. For example, an English KG usually contains more facts about the English-speaking society, while a Chinese KG contains more facts about the Chinese-speaking society. Considering that more KGs together can provide more comprehensive structural facts from various aspects, researchers have proposed many methods (Cf. surveys [10, 11, 12]) to fuse a pair of KGs into a unified one. The general process is to first identify equivalent entities between the candidate KGs, and then let them serve as the “bridge entities” to link the candidate KGs into a unified one.
@@ I Introduction
However, all the existing EA methods are specifically designed for aligning only a pair of candidate KGs, which can cover limited knowledge for downstream applications. To the best of our knowledge, none of the existing EA methods consider aligning multiple (more than two) KGs. For example, Fig. 1 shows a case where four candidate KGs need to be aligned, which triggers a more challenging research problem.
There are three obstacles that prevent traditional pair-wise KG alignment methods from being applied to the multiple KG alignment problem studied in this work. First, existing EA datasets are particularly constructed for pair-wise KG alignment. At present, there is no benchmark dataset that can support the study of aligning multiple KGs. Second, if we trivially adapt existing methods to this task, they need to be separately executed multiple times since they can only utilize the local (pair-wise) alignment information in a single pass, which is inefficient. Last and most importantly, they cannot capture some useful global (beyond pair-wise) alignment information, affecting the final alignment performance. For the example in Fig. 1, when they align KG-en with KG-fr, and align KG-en with KG-zh independently, the entities of the three KGs will be projected into four separate embedding spaces. Consequently, it is easy to yield inconsistent results like $e_{1}^{(en)}\equiv e_{1}^{(fr)}$, $e_{1}^{(en)}\equiv e_{1}^{(zh)}$, and $e_{1}^{(fr)}\equiv e_{2}^{(zh)}$ (wrong), violating the transitivity of the alignment relationship.
To fill the research gap in this field, in this work, we innovatively study the problem of aligning multiple KGs in a single pass. First of all, to facilitate the study of this problem, we construct two new benchmark datasets that contain multiple KGs to be aligned. Then, we formulate a novel research problem of aligning multiple KGs concurrently. Finally, we propose an effective framework named MultiEA to solve the problem. As illustrated in Fig. 1, our MultiEA can concurrently align the four candidate KGs in a single pass, which is more practical. In addition, it embeds the entities of all the candidate KGs into a unified space and thus can capture useful global alignment information, which helps achieve better alignment performance. Recall the example above, if MultiEA pulls $e_{1}^{(en)}$ and $e_{1}^{(fr)}$ closer and pulls $e_{1}^{(en)}$ and $e_{1}^{(zh)}$ closer. Then, $e_{1}^{(fr)}$ and $e_{1}^{(zh)}$ will definitely be pulled closer as well, leading to the consistent result: $e_{1}^{(en)}\equiv e_{1}^{(fr)}$, $e_{1}^{(en)}\equiv e_{1}^{(zh)}$, and $e_{1}^{(fr)}\equiv e_{1}^{(zh)}$ (right).
@@
  • We conduct extensive experiments on the two constructed benchmark datasets. The results demonstrate that MultiEA can effectively and efficiently align multiple KGs in a single pass. We release our source codes to facilitate further research on this problem.
II Related Work
@@
Trans-based EA Methods. Early EA methods exploit KG embedding methods [35], e.g., TransE [19], to project a pair of candidate KGs into a common low-dimensional Euclidean space by letting the embeddings of head entity $h$, relation $r$, and tail entity $h$ satisfy the triangular relationship: $h+r\approx t$. Given a set of pre-aligned entity pairs (i.e., seed alignment labels), they minimize the distance between these equivalent entity pairs in the embedding space. Thus, more potential equivalent entity pairs can be discovered after model optimization. Representative Trans-based EA methods include JEwP [13], JAPE [16], MTransE [15], IPTransE [14], BootEA [17], TransEdge [36], MultiKE [18], etc.
GNN-based EA Methods. Motivated by the extraordinary success of graph neural networks (GNNs) in extracting the structural features of graphs, GCN-Align [20] first introduces GCN [34] into EA, which has shown superior performance than Trans-based EA Methods, e.g., JEwP [13], MTransE [15] and JAPE [16]. Further, AVR-GCN [21] deftly combines the graph convolutional operator in GNNs and the translation operator in Trans-based methods to extract better entity embeddings. Afterward, a series of following GNN-based EA methods, e.g.,  [22, 23, 24, 25, 26, 27, 30, 31, 33, 32, 28, 29, 37, 38, 39, 40] are proposed. Recently, RREA [28] insightfully summarizes KG embedding-based EA methods and GNN-based EA methods into a unified framework consisting of two modules: shape-builder and alignment. The former constraints KGs into a specific distribution in the embedding space, and the latter minimizes the distance between the embeddings of pre-aligned entity pairs.
@@ III Preliminaries
As shown in Fig. 1, KG-en, KG-zh, KG-fr, and KG-ja are four candidate KGs that need to be aligned, where the four entities marked with red dashed lines refer to the same object, i.e., “Sweden”. We treat the four equivalent entities as one of the seed alignment labels, to help our EA algorithm find more potential equivalent entities. Finally, we can link (merge) the four KGs based on these detected equivalent entities across the four KGs.
IV Framework
@@ IV-B Training for Alignment
In this work, we propose three strategies for minimizing the distance among equivalent entities. The ideas are intuitively shown in Fig. 2. For a positive example $l=(e^{(1)},e^{(2)},...,e^{(m)},...,e^{(M)})\in\mathcal{S}$, we firstly obtain its associated entity embeddings from the KG encoder and denote them as: $(\mathbf{h}^{(1)},\mathbf{h}^{(2)},...,\mathbf{h}^{(m)},...,\mathbf{h}^{(M)})$. Then, the distance $d(l)$ is minimized by three strategies as follows.
(1) Moving toward Mean. This strategy lets all the equivalent entities approach their mean, as illustrated in Fig. 2(a). We first compute the mean as follows:
@@
(2) Moving toward Anchor. In practice, the candidate KGs are usually unbalanced due to various realistic factors. For instance, on our constructed multi-lingual KG dataset DBP-4, the English KG contains many more triples than the other three KGs (Cf. Table I). Therefore, as shown in Fig. 2(b), we designate entities from KG DBP-4: en as anchor entities and let the equivalent entities from the other KGs approach these anchor entities. This idea is formally described as follows:
@@
(3) Moving toward Each Other. This strategy lets equivalent entities approach each other. As illustrated in Fig. 2(c), we minimize the distances between all possible pairs of equivalent entities, formally described as follows:
@@
Input: $M$ KGs $\{\mathcal{G}^{(1)},\mathcal{G}^{(2)},...,\mathcal{G}^{(m)},...,\mathcal{G}^{(M)}\}$, and labels $\mathcal{S}$.
1:  Randomly initialize model parameters;
2:  Select one distance metric function from Eqs. (11, 12), Eq. (13), or Eq. (14);
3:  Augment KG data by Eqs. (1, 2);
4:  Get neighbor relation-entity tuples by Eq. (3);
5:  while not converge do
6:     Compute projection matrices by Eq. (5);
7:     Compute attention coefficients by Eqs. (7, 8);
8:     Perform multiple layers of aggregation by Eq. (4);
9:     Randomly sample negative examples by Eq. (10);
10:     Compute loss by Eq. (9);
11:     Update model parameters by gradient descent;
12:  end while
IV-C Training Time Complexity
@@
In the KG encoding phase (Cf. Section IV-A), the time complexity involves three main computational operations. First, the aggregation operation described by Eq. (4) has time complexity ${\mathcal{O}}(|\mathcal{E}|\cdot d^{2})$. Second, the computation of relation-specific projection matrices described by Eq. (5) has time complexity ${\mathcal{O}}(|\mathcal{R}|\cdot d^{2})$. Finally, as described by Eqs. (7, 8), the computation of attention coefficients has time complexity ${\mathcal{O}}(|\mathcal{T}|\cdot(d+d+d^{2}))$. Considering that in practice, the embedding dimensionality $d$ is a relatively small value, the time complexity of KG encoding is ${\mathcal{O}}(|\mathcal{E}|+|\mathcal{R}|+|\mathcal{T}|)$.
In the alignment phase (Cf. Section IV-B), as described by Eq. (12) and Eq. (13), both the mean strategy and the anchor strategy require time complexity ${\mathcal{O}}(M)$. As described by Eq. (14), the each other strategy requires time complexity ${\mathcal{O}}(M^{2})$. As described by Eq. (9), taking the number of labels into consideration. The mean strategy and the anchor strategy have time complexity of ${\mathcal{O}}(|\mathcal{S}|\cdot\eta\cdot M^{2})$, and each other strategy has time complexity ${\mathcal{O}}(|\mathcal{S}|\cdot\eta\cdot M^{3})$.
@@ V-A Benchmark Datasets
V-B Evaluation Metric
@@
TABLE II: Comparison of multiple KG alignment accuracy. The best results are bolded, and the second best results are underlined. The notation denotes that we re-implement the baseline methods to make them compatible with the task.
Methods DBP-4 DWY-3
MTransE 3.40% 19.72% 32.64% 25.83% 55.94% 65.31%
IPTransE 4.11% 23.36% 39.55% 28.74% 60.02% 68.84%
GCN-Align 4.48% 27.62% 44.94% 32.46% 63.18% 73.91%
KECG 5.62% 33.35% 49.61% 34.08% 67.51% 77.52%
NAEA 6.14% 34.13% 48.07% 35.17% 69.88% 79.43%
PSR 7.84% 38.17% 50.97% 37.55% 71.85% 80.13%
MultiEA+mean-infer 5.95% 40.48% 51.02% 36.68% 69.33% 79.65%
MultiEA+anchor-infer 6.32% 40.33% 52.08% 37.54% 70.21% 79.19%
MultiEA+each-infer 6.17% 41.70% 52.00% 36.76% 71.52% 79.18%
MultiEA+mean+infer 10.03% 48.18% 59.89% 40.54% 73.90% 81.24%
MultiEA+anchor+infer 10.58% 48.80% 59.95% 42.33% 74.13% 83.54%
MultiEA+each+infer 10.25% 50.38% 61.75% 41.19% 75.44% 85.42%
V-C Baselines and Variants
@@
Recall that in Section IV-B, we introduce three different alignment strategies. Here, we use +mean, +anchor, and +each to denote the three strategies, respectively. In Section IV-D, we propose the inference enhancement module. Here, we use +infer or -infer to indicate whether the module is equipped. By considering all the permutations, we set the following six variants for our MultiEA:
@@ V-D Implementation Details
Figure 5: Visualization of the embeddings of the ten groups of equivalent entities listed in Fig. 3 and Fig. 4.
@@ V-E Multiple KG Alignment
V-F Case Study
To intuitively show the effectiveness of our MultiEA, in Fig. 3 and Fig. 4, we list ten groups of equivalent entities discovered by MultiEA (the best-performing variant MultiEA+each+infer) on the two datasets. As we can see, MultiEA can effectively discover equivalent entities across all the candidate KGs.
Further, we visualize the embeddings of these entities. Specifically, we utilize the t-SNE algorithm [58] to project them into the 2-dimensional Euclidean space. Fig. 5(a) and Fig. 5(b) show the results on DBP-4 and DWY-3, respectively, where different colors and different shapes mark the entities from different KGs, and the equivalent entities are connected by black lines. As we can see, on both datasets, the ten groups of equivalent entities are obviously gathered together, indicating the effectiveness of the alignment mechanism of MultiEA. On the other hand, a proper margin is maintained between different groups, avoiding the risk of model overfitting. The result on DWY-3 shows a relatively better distribution. This is consistent with the higher accuracy on DWY-3 as reported in Table II.
V-G Efficiency Study
In this experiment, we demonstrate the efficiency advantage of our MultiEA in aligning multiple KGs. As shown in Table III and Table IV, we first report the experimental results of our re-implemented GNN-based EA baselines as described in Section V-E. That is, they align multiple candidate KGs by splitting this task into multiple separate pair-wise sub-tasks, as marked by in the tables. Then, we apply the “each other” alignment strategy and the inference enhancement module of our MultiEA+each+infer to each baseline X, as denoted as MultiEA (X) in the tables.